A wave 1.30 m long travels 80.0 m in 25 s. What is the period of the wave?


1
Expert's answer
2021-06-16T15:38:27-0400

"T=\frac{1}{\nu}=\frac{\lambda}{v}=\frac{\lambda}{\frac lt}=\frac{\lambda t}{l}=\frac{1.3\cdot 25}{80}=406~ms."